### v1

# Subspace Ordering for Maximum Response Preservation in Sufficient Dimension Reduction

Abstract
Sufficient dimension reduction (SDR) methods aim to identify a dimension reduction subspace (DRS) that preserves all the information about the conditional distribution of a response given its predictor. Traditional SDR methods determine the DRS by solving a method-specific generalized eigenvalue problem and selecting the eigenvectors corresponding to the largest eigenvalues. In this article, we argue against the long-standing convention of using eigenvalues as the measure of subspace importance and propose alternative ordering criteria that directly assess the predictive relevance of each subspace. For a binary response, we introduce a subspace ordering criterion based on the absolute value of the independent Student’s T-statistic. Theoretically, our criterion identifies subspaces that achieve the local minimum Bayes’ error rate and yields consistent ordering of directions under mild regularity conditions. Additionally, we employ an F-statistic to provide a framework that unifies categorical and continuous responses under a single subspace criterion. We evaluate our proposed criteria within multiple SDR methods through extensive simulation studies and applications to real-data. Our empirical results demonstrate the efficacy of reordering subspaces using our proposed criteria, which generally improves classification accuracy and subspace estimation compared to ordering by eigenvalues.
Keywords: Central subspace, Eigenvalues, Selection criteria, Spectral decomposition, Supervised learning
1 Introduction
The dimension reduction subspace (DRS) in sufficient dimension reduction (SDR) is often composed of a subset of the leading eigenvectors of a method-specific generalized eigenvalue problem. That is, the DRS is spanned by the eigenvectors corresponding to the largest eigenvalues in magnitude. This approach to choosing eigenvectors based upon eigenvalues is common practice and has the intent of maximizing the variability of the data in the chosen subspace. However, for supervised statistical learning, we argue that the use of eigenvalues to determine a relevant DRS is generally flawed. This problem results because maximum variability does not guarantee that the selected subspace best preserves the relationship between the predictors and the response, which is the primary goal of SDR.
Let $Y$ be the response of the predictor $\mathbf{X}\in\mathbb{R}^{p}$. Then, the goal of SDR is to project $\mathbf{X}$ onto the smallest possible subspace $\mathcal{S}\subseteq\mathbb{R}^{p}$ without any loss of information with respect to $Y|\mathbf{X}$. In SDR, the dimension reduction is usually constrained to a linear transformation $\mathbf{P}_{\mathcal{S}}\mathbf{X}$, where $\mathbf{P}_{\mathcal{S}}\in\mathbb{R}^{p\times p}$ is the projection matrix onto $\mathcal{S}$ in the standard inner product. Let $\sim$” and $\perp\!\!\!\perp$” denote “distributed as” and “independent of,” respectively. We formally define the dimension reduction as a sufficient linear reduction if it satisfies at least one of the following: (i) $Y\perp\!\!\!\perp\mathbf{X}|\mathbf{P}_{\mathcal{S}}\mathbf{X}$, (ii) $\mathbf{X}|\left(Y,\mathbf{P}_{\mathcal{S}}\mathbf{X}\right)\sim\mathbf{X}|\mathbf{P}_{\mathcal{S}}\mathbf{X}$, or (iii) $Y|\mathbf{X}\sim Y|\mathbf{P}_{\mathcal{S}}\mathbf{X}$. Thus, $\mathcal{S}$ is called a DRS and is said to be a minimum DRS for $Y|\mathbf{X}$ if $\text{dim}(\mathcal{S})\leq\text{dim}(\mathcal{S}_{DRS})$, where $\text{dim}(\cdot)$ denotes dimension and $\mathcal{S}_{DRS}$ represents all other DRS. The minimum DRS, however, may not be unique (e.g., see Section 6.3 of Cook, (1998)). To address this issue, Cook, (1998) introduced the central subspace (CS), defined as the intersection of all DRS, and denoted it as $\mathcal{S}_{Y|\mathbf{X}}=\cap\mathcal{S}_{DRS}$. Given that $\cap\mathcal{S}_{DRS}$ itself is a DRS and, under mild conditions given by Cook, (1998), $\mathcal{S}_{Y|\mathbf{X}}$ exists and is the unique minimum DRS. Thus, most SDR methods estimate the CS, or at least a portion of the CS, as the target subspace. Throughout the paper, we assume the existence of the CS. Let $d<p$ be the structural dimension of the CS, and let $\boldsymbol{\beta}=\left(\boldsymbol{\beta}_{1},\ldots,\boldsymbol{\beta}_{d}\right)\in\mathbb{R}^{p\times d}$ be a basis matrix of $\mathcal{S}_{Y|\mathbf{X}}$. Thus, SDR methods can reduce the dimensionality of predictors to $\boldsymbol{\beta}^{\top}\mathbf{X}\in\mathbb{R}^{d}$ for subsequent supervised learning without loss of information. For a comprehensive review of SDR, see Li, (2018).
Most classical SDR methods rely on slicing the data into $H$ contiguous non-overlapping intervals to construct functions of the first two conditional moments with the goal of recovering the CS. Notable examples of this approach include sliced inverse regression (SIR), introduced by Li, (1991), and sliced average variance estimation (SAVE), proposed by Cook and Weisberg, (1991). This slicing-based framework is particularly suited to a continuous response, where slicing facilitates tractable estimation of conditional expectations and yields a DRS that is contained in $\mathcal{S}_{Y|\mathbf{X}}$ for sufficiently large $H$. When the response is categorical, however, slicing becomes trivial since the grouping of observations is explicit by the $H$ distinct populations. Moreover, with a categorical response, the emphasis is often on optimizing a classification rule rather than examining other aspects of $Y|\mathbf{X}$. Thus, Cook and Yin, (2001) proposed the central discriminant subspace (CDS) as an alternative target subspace to the CS for discriminant analysis. Let the Bayes’ rule be $\phi(\mathbf{X})\coloneqq\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\mathbf{X}\right)$. For a subspace $\mathcal{S}\subseteq\mathbb{R}^{p}$, let $\phi_{\mathcal{S}}(\mathbf{X})\coloneqq\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\mathbf{P}_{\mathcal{S}}\mathbf{X}\right)$. Moreover, for any basis matrix $\boldsymbol{\beta}$ such that $\text{span}(\boldsymbol{\beta})=\mathcal{S}$, we have $\phi_{\mathcal{S}}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\Pr\left(Y=h|\boldsymbol{\beta}^{\top}\mathbf{X}\right)$. Thus, if $\mathcal{S}$ satisfies $\phi_{\mathcal{S}}(\mathbf{X})=\phi(\mathbf{X})$, then it is a discriminant subspace. The CDS, denoted as $\mathcal{S}_{D(Y|\mathbf{X})}\subseteq\mathcal{S}_{Y|\mathbf{X}}$, is then defined as the intersection of all discriminant subspaces, given that the intersection itself is a discriminant subspace.
Regardless of whether the emphasis is on the CS or CDS, most SDR methods rely on eigenvectors to estimate $\text{span}(\boldsymbol{\beta})$. These eigenvectors are typically ordered by the magnitude of their associated eigenvalues with the assumption that subspaces corresponding to relatively large eigenvalues are more informative. However, large eigenvalues generally only represent larger data variability in the subspace and do not guarantee that the predictor-response relationship is preserved (e.g., see Huber, (1985)). Thus, we propose new and more appropriate subspace ordering criteria that explicitly capture the predictive information in each direction, thereby ensuring that the leading subspaces align with the intended goal of the supervised learner.
More specifically, when the response is binary, we propose using the absolute value of the independent Student’s T-statistic, introduced by Welch, (1947), as a simple yet effective importance measure for a DRS. We naturally extend the idea of a subspace ordering criterion to a categorical response with more than two populations by employing an F-statistic, introduced by Fisher, (1925). Furthermore, within the slicing-based framework, we demonstrate that an F-statistic can also serve as a subspace criterion in the continuous response setting. That is, by emphasizing maximum slice separation, we can often identify subspaces that best capture the conditional moments. Although methods exist to determine the structural dimension $d$ of the CS (e.g., see Chapters 9 and 10 in Li, (2018) and see Zeng et al., (2024)), to the best of our knowledge, we are the first to propose reordering the DRS by a criterion different from the eigenvalue magnitudes. We believe that, while the intuition of these proposed criteria is simple, its simplicity offers a novel and interpretable importance measure for subspaces that addresses a previously overlooked but crucial aspect of subspace selection—namely, the need to directly assess predictive relevance rather than defaulting to eigenvalue magnitude.
The remainder of the paper proceeds as follows. In Section 2, we establish the notation used throughout the paper and provide a brief review of relevant methodologies. In Section 3, we provide a simple example to illustrate the potential gain of reordering a DRS by an Student’s T-statistic rather than the eigenvalue magnitude. We then study the theoretical properties of the criterion to establish it as a consistent and relevant importance measure for a DRS. In Section 4, we establish a criterion using an F-statistic for a categorical or continuous response. In Sections 5 and 6, we present the simulation studies and real-data applications, respectively. In Section 7, we discuss additional work and conclude our findings.
2 A Review of Methodologies
2.1 Notation
First, we define the following notation that will be used throughout the paper. Let $\mathbb{R}^{m\times n}$ represent the set of all $m\times n$ real matrices, $\mathbb{S}^{p}\subset\mathbb{R}^{p\times p}$ represent the set of $p\times p$ real symmetric matrices, and $\mathbb{S}^{p}_{+}\subset\mathbb{S}^{p}$ denote the interior of the cone of $p\times p$ real symmetric positive-definite matrices. Let $\oplus$ denote the direct sum such that $\mathbf{A}\oplus\mathbf{B}=\begin{bmatrix}\mathbf{A}&\mathbf{0}\\ \mathbf{0}&\mathbf{B}\end{bmatrix}$. For $h=1,\ldots,H$, let $\Pi_{h}$ represent the $h^{\text{th}}$ distinct interval of $Y$ with the a priori membership $\pi_{h}=\Pr(Y=h)$, where $\sum_{h=1}^{H}\pi_{h}=1$. Let $\boldsymbol{\Sigma}_{h}\in\mathbb{S}^{p}_{+}$ denote the population covariance matrix for $\Pi_{h}$, and let $\boldsymbol{\mu}_{h}\in\mathbb{R}^{p\times 1}$ be the population mean vector for $\Pi_{h}$.
2.2 Discriminant Analysis
In Section 1, we defined the CDS as the intersection of all subspaces that preserves Bayes’ rule, $\phi(\mathbf{X})=\arg\max_{h=1,\ldots,H}\Pr(Y=h\mid\mathbf{X})$. Because the conditional class probability $\Pr(Y=h\mid\mathbf{X})$ is generally unknown and lacks a closed-form expression, a common approach is to consider settings in which it is tractable. One such setting is the multivariate normal population model for the predictor $\mathbf{X}\in\mathbb{R}^{p}$ given the categorical response $Y\in\{1,\ldots,H\}$, $H\geq 2$, which is given by
$\mathbf{X}\mid\{Y=h\}\sim\mathcal{N}(\boldsymbol{\mu}_{h},\boldsymbol{\Sigma}_{h}),\quad h=1,\ldots,H.$ (1)
If no assumption is made on $\boldsymbol{\Sigma}_{1},\ldots,\boldsymbol{\Sigma}_{H}$, then, under model (1), the Bayes’ rule for classification is
$\phi_{QDA}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\left\{\log\pi_{h}+\frac{1}{2}\log|\boldsymbol{\Sigma}_{h}|-\frac{1}{2}(\mathbf{X}-\boldsymbol{\mu}_{h})^{\top}\boldsymbol{\Sigma}_{h}^{-1}(\mathbf{X}-\boldsymbol{\mu}_{h})\right\},$ (2)
which is commonly referred to as the Bayes’ quadratic discriminant analysis (QDA) rule.
Because the QDA rule does not require homoscedasticity, the Bayes’ rule for classification is a quadratic function of $\mathbf{X}$. Although expression (2) gives the Bayes’ rule in its standard form, we can algebraically reformulate it to explicitly isolate the linear and quadratic components by subtracting the discriminant function for a reference class (e.g., class 1),
$\phi_{QDA}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\left\{c_{h}-\mathbf{X}^{\top}(\boldsymbol{\Sigma}_{h}^{-1}\boldsymbol{\mu}_{h}-\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\mu}_{1})+\frac{1}{2}\mathbf{X}^{\top}(\boldsymbol{\Sigma}_{h}^{-1}-\boldsymbol{\Sigma}_{1}^{-1})\mathbf{X}\right\},$ (3)
where $c_{h}=\log\pi_{h}+\log|\boldsymbol{\Sigma}_{h}|/2+\boldsymbol{\mu}_{h}^{\top}\boldsymbol{\Sigma}_{h}^{-1}\boldsymbol{\mu}_{h}/2$ is a constant term that does not depend on $\mathbf{X}$. Thus, expression (3) yields that optimal dimension reduction under the QDA model is characterized by the subspace
$\mathcal{L}\coloneqq\text{span}\{\boldsymbol{\Sigma}_{h}^{-1}\boldsymbol{\mu}_{h}-\boldsymbol{\Sigma}_{1}^{-1}\boldsymbol{\mu}_{1}\mid h=2,\ldots,H\}\subseteq\mathbb{R}^{p},$ (4)
which corresponds to the linear components, and
$\mathcal{Q}\coloneqq\text{span}\{\boldsymbol{\Sigma}_{h}^{-1}-\boldsymbol{\Sigma}_{1}^{-1}\mid h=2,\ldots,H\}\subseteq\mathbb{R}^{p},$ (5)
which corresponds to the quadratic components. This result is formalized in the following lemma. All proofs for lemmas and theorems are given in the Supplementary Material.
Lemma 1.
Let $\mathcal{L}$ and $\mathcal{Q}$ be as defined in (4) and (5), respectively. Then, under model (1), $\mathcal{S}_{Y\mid\mathbf{X}}=\mathcal{S}_{D(Y\mid\mathbf{X})}=\mathcal{L}\cup\mathcal{Q}$.
If we assume $\boldsymbol{\Sigma}_{1}=\ldots=\boldsymbol{\Sigma}_{H}=\boldsymbol{\Sigma}$, then the Bayes’ rule simplifies to what is commonly referred to as the Bayes’ linear discriminant analysis (LDA) rule, which reduces (3) to
$\phi_{LDA}(\mathbf{X})=\arg\max_{h=1,\ldots,H}\left\{(\boldsymbol{\mu}_{h}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}^{-1}(\mathbf{X}-[\boldsymbol{\mu}_{h}-\boldsymbol{\mu}_{1}]/2)\right\}.$ (6)
To distinguish the $H$ populations under such a linear model, we require at most H1H-1 directions. Thus, we consider the subspace
$\mathcal{B}\coloneqq\text{span}\{\boldsymbol{\Sigma}^{-1}(\boldsymbol{\mu}_{h}-\boldsymbol{\mu}_{1})\mid h=2,\ldots,H\}\subseteq\mathbb{R}^{p}.$ (7)
Then, under the LDA rule, the following lemma establishes that $\mathcal{B}$ coincides with both the CS and CDS.
Lemma 2.
Let $\mathcal{B}$ be as defined in (7) and $\boldsymbol{\Sigma}_{1}=\ldots=\boldsymbol{\Sigma}_{H}=\boldsymbol{\Sigma}$. Then, under model (1), $\mathcal{S}_{Y\mid\mathbf{X}}=\mathcal{S}_{D(Y\mid\mathbf{X})}=\mathcal{B}$.
Both the QDA and LDA discriminant subspaces contain a linear component. However, the linear component $\mathcal{L}$ of the discriminant subspace under the QDA model generally differs from the linear component $\mathcal{B}$ in LDA. Although $\mathcal{L}\neq\mathcal{B}$, the subspaces coincide when the heteroscedastic covariance matrices in the QDA model are pooled according to $\boldsymbol{\Sigma}=\sum_{h=1}^{H}\pi_{h}\boldsymbol{\Sigma}_{h}$. In that case, Zhang and Mai, (2019) showed that $\mathcal{L}$ and $\mathcal{B}$ span the same subspace when combined with $\mathcal{Q}$, which yields the following identity.
Lemma 3.
Let $\mathcal{L}$, $\mathcal{Q}$, and $\mathcal{B}$ be as defined in (4), (5), and (7), respectively. Then, under model (1), $\mathcal{S}_{Y\mid\mathbf{X}}=\mathcal{S}_{D(Y\mid\mathbf{X})}=\mathcal{L}\cup\mathcal{Q}=\mathcal{B}\cup\mathcal{Q}.$
Here, $\phi_{\text{QDA}}$ and $\phi_{\text{LDA}}$ characterize supervised classification decision rules that assign an unlabeled observation $\mathbf{X}\in\mathbb{R}^{p}$ to a distinct population. While the subspace structure of these rules governs the directions along which class separation occurs, their effectiveness is ultimately quantified by the associated Bayes’ error rate, defined as the probability of misclassification under the optimal rule. For example, if we consider the simple case where $H=2$ such that $\pi_{1}=\pi_{2}=1/2$, then we can easily show (e.g., Johnson and Wichern, (2007)) that the optimal error rate (OER) under the LDA decision rule is given by
$\Phi\left(-\frac{1}{2}\sqrt{(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})^{\top}\boldsymbol{\Sigma}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})}\right),$ (8)
where $\Phi(\cdot)$ is the cumulative distribution function (CDF) of a standard normal random variable. This expression yields the optimal Bayes’ error rate based on the one-dimensional (1D) projection of $\mathbf{X}\in\mathbb{R}^{p}$ onto the direction $\boldsymbol{\Sigma}^{-1}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})$, which spans $\mathcal{B}$. Thus, by Lemma 2, we have $\mathcal{B}=\mathcal{S}_{D(Y\mid\mathbf{X})}$ under the LDA rule, which confirms that maximum population separation, and consequently the OER, is achieved when we project observations onto the CDS. Therefore, for any subspace spanned by an arbitrary nonzero vector $\mathbf{v}\in\mathbb{R}^{p}$, the Bayes’ error rate in that subspace will exceed the global OER in (8) unless $\mathbf{v}$ spans the same subspace as $\mathcal{B}$. However, among a set of $p$ candidate directions $\mathbf{v}_{1},\ldots,\mathbf{v}_{p}$, we can use the Bayes’ OER to determine the subspace most aligned with $\mathcal{S}_{D(Y\mid\mathbf{X})}$. Thus, under model (1), we have $\mathbf{v}^{\top}_{j}\mathbf{X}|\{Y=h\}\sim\mathcal{N}(\mathbf{v}_{j}^{\top}\boldsymbol{\mu}_{h},\mathbf{v}^{\top}_{j}\boldsymbol{\Sigma}_{h}\mathbf{v}_{j})$, and, consequently, the 1D Bayes’ OER in the subspace defined by $\text{span}(\mathbf{v}_{j})$ is given by
$\varphi_{j}\coloneqq\begin{cases}\Phi\left(-\dfrac{1}{2}\dfrac{|\mu_{2j}^{\ast}-\mu_{1j}^{\ast}|}{\sigma^{\ast}_{j}}\right),&\sigma_{1j}^{\ast}=\sigma_{2j}^{\ast}=\sigma^{\ast}_{j}\\ \dfrac{1}{2}+\dfrac{1}{2}\Phi\left(\dfrac{\sigma_{1j}^{\ast}(\mu_{2j}^{\ast}-\mu_{1j}^{\ast})-\sigma_{2j}^{\ast}\tau}{{\sigma_{2j}^{2}}^{\ast}-{\sigma_{1j}^{2}}^{\ast}}\right)-\dfrac{1}{2}\Phi\left(\dfrac{\sigma_{1j}^{\ast}(\mu_{2j}^{\ast}-\mu_{1j}^{\ast})+\sigma_{2j}^{\ast}\tau}{{\sigma_{2j}^{2}}^{\ast}-{\sigma_{1j}^{2}}^{\ast}}\right)\\ \quad+\dfrac{1}{2}\Phi\left(\dfrac{\sigma_{2j}^{\ast}(\mu_{2j}^{\ast}-\mu_{1j}^{\ast})+\sigma_{1j}^{\ast}\tau}{{\sigma_{2j}^{2}}^{\ast}-{\sigma_{1j}^{2}}^{\ast}}\right)-\dfrac{1}{2}\Phi\left(\dfrac{\sigma_{2j}^{\ast}(\mu_{2j}^{\ast}-\mu_{1j}^{\ast})-\sigma_{1j}^{\ast}\tau}{{\sigma_{2j}^{2}}^{\ast}-{\sigma_{1j}^{2}}^{\ast}}\right),&\sigma_{2j}^{\ast}>\sigma_{1j}^{\ast},\end{cases}$ (9)
where $\mu_{hj}^{\ast}\coloneqq\mathbf{v}^{\top}_{j}\boldsymbol{\mu}_{h}$, ${\sigma_{hj}^{2}}^{\ast}\coloneqq\mathbf{v}^{\top}_{j}\boldsymbol{\Sigma}_{h}\mathbf{v}_{j}$, and $\tau\coloneqq\sqrt{(\mu_{2j}^{\ast}-\mu_{1j}^{\ast})^{2}+({\sigma_{2j}^{2}}^{\ast}-{\sigma_{1j}^{2}}^{\ast})\log\left({\sigma_{2j}^{2}}^{\ast}/{\sigma_{2j}^{2}}^{\ast}\right)}$. When $\sigma_{1j}^{\ast}=\sigma_{2j}^{\ast}$, (9) directly follows from (8), and when $\sigma_{1j}^{\ast}\neq\sigma_{2j}^{\ast}$, we use the corresponding expression derived by Wu and Hao, (2022). Note that in (8) and (9), we assume equal prior class probabilities for simplicity of presentation. However, this assumption is not essential, and all subsequent theoretical results remain valid with minor modifications.
2.3 Select SDR Methods
Table 1: Generalized eigenvalue formulations for select SDR methods.
Method $\mathbf{M}$ $\mathbf{N}$
PCA $\boldsymbol{\Sigma}_{\mathbf{X}}$ $\mathbf{I}_{p}$
SIR $\text{Cov}\left(\mathbb{E}[\mathbf{X}-\mathbb{E}\{\mathbf{X}|Y\}]\right)$ $\boldsymbol{\Sigma}_{\mathbf{X}}$
SAVE $\Sigma_{\mathbf{X}}^{1/2}\mathbb{E}\left[\left\{\mathbf{I}_{p}-\text{Cov}(\mathbf{Z}\mid Y)\right\}^{2}\right]\Sigma_{\mathbf{X}}^{1/2}$ $\boldsymbol{\Sigma}_{\mathbf{X}}$
SIR-II $\mathbb{E}\left[\text{Var}(\mathbf{Z}|Y)-\mathbb{E}\{\text{Var}(\mathbf{Z}|Y)\}\right]^{2}$ $\boldsymbol{\Sigma}_{\mathbf{X}}$
DR $\big\{\frac{1}{2}[2\mathbb{E}[\mathbb{E}^{2}(\mathbf{ZZ}^{T}\mid Y)]+2\mathbb{E}^{2}[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(Z^{T}\mid Y)]+$ $\boldsymbol{\Sigma}_{\mathbf{X}}$
$2\mathbb{E}\left[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(\mathbf{Z}\mid Y)\right]\mathbb{E}\left[\mathbb{E}(\mathbf{Z}\mid Y)\mathbb{E}(\mathbf{Z}^{T}\mid Y)\right]-2\mathbf{I}_{p}]\}^{1/2}$
SSDR $\left(\mathcal{L},\boldsymbol{\Sigma}_{2}-\boldsymbol{\Sigma}_{1},\ldots,\boldsymbol{\Sigma}_{H}-\boldsymbol{\Sigma}_{1}\right)$ $\left(\mathcal{L},\boldsymbol{\Sigma}_{2}-\boldsymbol{\Sigma}_{1},\ldots,\boldsymbol{\Sigma}_{H}-\boldsymbol{\Sigma}_{1}\right)^{\top}$ $\mathbf{I}_{p}$
Li, (2007) has shown that most SDR methods can be formulated as a generalized eigenvalue problem given by
$\mathbf{M}\mathbf{v}_{j}=\lambda_{j}\mathbf{N}\mathbf{v}_{j},j=1,\ldots,p,$
where $\mathbf{M}\in\mathbb{R}^{p\times p}$ is a method-specific symmetric kernel matrix, and $\mathbf{N}\in\mathbb{S}^{p}_{+}$ is often taken to be the common covariance matrix, denoted as $\boldsymbol{\Sigma}_{\mathbf{X}}$. Additionally, $\mathbf{v}_{1},\ldots,\mathbf{v}_{p}$ are the eigenvectors satisfying $\mathbf{v}_{j}^{\top}\mathbf{N}\mathbf{v}_{\ell}=1$ for $j=\ell$ and $\mathbf{v}_{j}^{\top}\mathbf{N}\mathbf{v}_{\ell}=0$ for $\ell\neq j$, corresponding to the eigenvalues $\lambda_{1}\geq\ldots\geq\lambda_{p}$. We denote the generalized eigenvalue problem with matrices $\mathbf{M}$ and $\mathbf{N}$ as $\text{GEV}(\mathbf{M},\mathbf{N})$. Let $\mathbf{Z}\coloneqq\boldsymbol{\Sigma}_{\mathbf{X}}^{-1/2}\{\mathbf{X}-\mathbb{E}(\mathbf{X})\}$. Then, we summarize the generalized eigenvalue problem for the following SDR methods in Table 1: principal components analysis (PCA) introduced by Pearson, (1901), SIR by Li, (1991), SAVE by Cook and Weisberg, (1991), a variant of SIR that considers second-order moments referred to as sliced inverse regression II (SIR-II) by Li, (1991), directional regression (DR) by Li and Wang, (2007), which aims to achieve an exhaustive estimate of the CS, and stabilized sufficient dimension reduction (SSDR) by Boonstra et al., (2025), which adjusts for heteroscedasticity and estimates a similar subspace as the CDS under QDA.
[…]
3 Subspace Ordering Criterion for Binary Response Preservation
3.1 Illustrative Example of Potential Gain by Reordering Subspaces
To assess the discriminant information of a DRS, we consider the independent Student’s T-statistic introduced by Welch, (1947),
$T\coloneqq\frac{\overline{x}_{2}-\overline{x}_{1}}{\sqrt{s^{2}_{2}/n_{1}+s^{2}_{1}/n_{2}}}.$ (10)
Fan and Fan, (2008) introduced the absolute value of (10) as a now widely used screening tool to select relevant features in high-dimensional data analysis and reduce computational burden (e.g., see Thudumu et al., (2020) and Fan et al., (2020)). In their theoretical justification of the screening criterion, Fan and Fan, (2008) showed that the absolute value of (10) for the $j^{\text{th}}$ feature converges in probability to $\left|\mu_{1j}-\mu_{2j}\right|/\sqrt{\sigma^{2}_{1j}/n_{1}+\sigma^{2}_{2j}/n_{2}}$ under bounded variance and minimal signal assumptions. Although they did not explicitly label this expression, the authors use it throughout their asymptotic arguments (e.g., Theorem 3), and it functions as a population signal-to-noise ratio (SNR) for predictors. Motivated by their work, we extend the notion of the population SNR to a DRS.
For each direction $\mathbf{v}_{j}$, we define the projected population SNR between populations as
$\Delta_{j}\coloneqq\frac{\left|\mathbf{v}^{\top}_{j}(\boldsymbol{\mu}_{2}-\boldsymbol{\mu}_{1})\right|}{\sqrt{\pi_{2}\mathbf{v}^{\top}_{j}\boldsymbol{\Sigma}_{2}\mathbf{v}_{j}+\pi_{1}\mathbf{v}^{\top}_{j}\boldsymbol{\Sigma}_{1}\mathbf{v}_{j}}},\quad j=1,\ldots,p.$ (11)
To estimate this quantity from data, we define the first two sample moments in the subspace spanned by $\widehat{\mathbf{v}}_{j}$ as
$\bar{x}^{*}_{hj}\coloneqq\frac{1}{n_{h}}\sum_{i=1}^{n_{h}}\widehat{\mathbf{v}}_{j}^{\top}\mathbf{x}_{hi},\quad s_{hj}^{2^{*}}\coloneqq\frac{1}{n_{h}-1}\sum_{i=1}^{n_{h}}\left(\widehat{\mathbf{v}}_{j}^{\top}\mathbf{x}_{hi}-\bar{x}^{*}_{hj}\right)^{2},$
where $\widehat{\pi}_{h}\coloneqq n_{h}/\sum_{h=1}^{H}n_{h}$. Thus, we define the sample analogue of the SNR in a DRS as
$T_{j}\coloneqq\frac{|\bar{x}^{\ast}_{2j}-\bar{x}^{\ast}_{1j}|}{\sqrt{\widehat{\pi}_{2}s_{2j}^{2\ast}+\widehat{\pi}_{1}s_{1j}^{2\ast}}},\quad j=1,\ldots,p.$ (12)
Here, $T_{j}$ serves as an estimate of $\Delta_{j}$ and quantifies the standardized separation between the two class means along the direction $\widehat{\mathbf{v}}_{j}$. Thus, we can use $T_{j}$ to evaluate and rank the components of a DRS.
To illustrate $T_{j}$ as a subspace criterion, we consider a simple example in which the leading eigenvectors do not necessarily correspond to the subspace that best preserves the relationship between the response and predictors. For this example, we focus on the simple SDR method of PCA. From Table 1, PCA can be formulated as solving $\text{GEV}(\boldsymbol{\Sigma}_{X},\mathbf{I}_{p})$, and the resulting DRS is defined by $\text{span}(\boldsymbol{\Sigma}_{X})$. Consider the covariance structure $\boldsymbol{\Sigma}_{X}=\text{diag}(3,2,1)$, a diagonal matrix with diagonal entries 3, 2, and 1. One can easily show that the eigenvectors of $\boldsymbol{\Sigma}_{X}$ are $\mathbf{v}_{1}=(1,0,0)^{\top}$, $\mathbf{v}_{2}=(0,1,0)^{\top}$, and $\mathbf{v}_{3}=(0,0,1)^{\top}$, and the respective eigenvalues are $\lambda_{1}=3$, $\lambda_{2}=2$, and $\lambda_{3}=1$. Thus, because $\lambda_{1}>\lambda_{2}>\lambda_{3}$, the basis for the traditional DRS is given by $\mathbf{V}=(\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3})$, with $\mathbf{v}_{1}$ corresponding to the first DRS with the largest variance. However, as we will demonstrate, $\mathbf{v}_{1}$ does not necessarily span the subspace that best captures the relationship between the response and the predictors. In particular, when we consider a binary response, we show that $T_{j}$ provides a better criterion for maximum population separation.
Figure 1: Simulated data from two multivariate normal populations described in Section 3.1 in the subspaces spanned by the leading eigenvectors $\mathbf{v}_{1}$, $\mathbf{v}_{2}$, and $\mathbf{v}_{3}$. Vertical jitter is added for visualization purposes.
Next, consider the simple parameter configuration $\boldsymbol{\mu}_{1}=(0,\varepsilon,\alpha)^{\top}$ and $\boldsymbol{\mu}_{2}=(0,0,0)^{\top}$, where $0<|\varepsilon|\ll|\alpha|$, and assume both populations share the common covariance matrix $\boldsymbol{\Sigma}_{X}=\text{diag}(3,2,1)$. Clearly, the discriminant information lies almost entirely in the third coordinate through $\alpha$ while the contributions from the first and second coordinates are negligible for a relatively small $\varepsilon$. Thus, only $\mathbf{v}_{3}=(0,0,1)^{\top}$, the eigenvector with the smallest eigenvalue, contains the dominant signal. This result is captured in the corresponding $\Delta_{j}$ values for each subspace spanned by $\mathbf{v}_{j}$ that simplify to $\Delta_{j}=\left|\mathbf{v}^{\top}_{j}\boldsymbol{\mu}_{1}\right|/\sqrt{\lambda_{j}}$ because $\lambda_{j}=\mathbf{v}_{j}^{\top}\boldsymbol{\Sigma}\mathbf{v}_{j}$ in the PCA setting. Hence, for $\mathbf{v}_{1}$, $\mathbf{v}_{2}$, and $\mathbf{v}_{3}$, we have $\Delta_{1}=0$, $\Delta_{2}=|\varepsilon|/\sqrt{2}$, and $\Delta_{3}=|\alpha|$, respectively. Since $|\alpha|>|\varepsilon|$, we have $\Delta_{3}>\Delta_{2}>\Delta_{1}$, and in the event that $\varepsilon\approx\alpha$, we still have $\Delta_{3}>\Delta_{2}$ because $\lambda_{2}>\lambda_{3}$. Thus, $\Delta_{j}$ indicates that $\mathbf{v}_{3}$, not $\mathbf{v}_{1}$, defines the most important subspace in terms of classification. Therefore, we should reorder the DRS as $\mathbf{V}=(\mathbf{v}_{3},\mathbf{v}_{2},\mathbf{v}_{1})$.
To visualize this example, we generated 1,000 observations from each population, in which we took $\alpha=5$ and $\varepsilon=2$ and assumed each population followed a multivariate normal distribution. In Figure 1, the simulated data is given in each subspace defined by $\mathbf{v}_{1}$, $\mathbf{v}_{2}$, and $\mathbf{v}_{3}$, respectively. Clearly, superior class separation was achieved in $\mathbf{v}_{3}$ compared to $\mathbf{v}_{2}$ or $\mathbf{v}_{1}$. Moreover, this ordering was preserved in the estimated basis $\widehat{\mathbf{V}}=(\widehat{\mathbf{v}}_{1},\widehat{\mathbf{v}}_{2},\widehat{\mathbf{v}}_{3})$, computed from the simulated data, where the sample principal components were determined to be $\widehat{\mathbf{v}}_{1}=(1.00,0.01,-0.01)^{\top}$, $\widehat{\mathbf{v}}_{2}=(-0.01,1.00,-0.01)^{\top}$, and $\widehat{\mathbf{v}}_{3}=(0.01,0.01,1.00)^{\top}$. The corresponding eigenvalues were $\widehat{\lambda}_{1}=3.23$, $\widehat{\lambda}_{2}=2.06$, and $\widehat{\lambda}_{3}=1.01$. Under a traditional eigenvalue or variance-based approach, $\widehat{\mathbf{v}}_{1}$ would be prioritized as the most informative subspace. However, the SNRs computed via $T_{j}$ tell a different story. We found $T_{1}=0.003$, $T_{2}=1.36$, and $T_{3}=4.97$, clearly indicating that $\widehat{\mathbf{v}}_{3}$ was the most discriminative subspace. Consistent with this finding, the estimated Bayes’ error rates under the LDA decision rule for the one-dimensional subspaces spanned by $\widehat{\mathbf{v}}_{1}$, $\widehat{\mathbf{v}}_{2}$, and $\widehat{\mathbf{v}}_{3}$ were 0.5040, 0.2560, and 0.0045, respectively. Thus, $\widehat{\mathbf{v}}_{3}$, the eigenvector corresponding to the smallest eigenvalue and traditionally considered the least informative under PCA, was in fact the most important subspace for discrimination. Therefore, using $T_{j}$ rather than eigenvalue magnitude as a subspace ordering criterion corrects a fundamental misalignment in SDR for supervised learning by prioritizing subspaces that preserve the predictor–response relationship rather than the subspaces that merely aim to maximize overall variability. We formalize this criterion in the next section with theoretical results.
3.2 Theoretical Properties
Here, we study the theoretical properties of $\Delta_{j}$ as an importance measure for a DRS by establishing its connection to the CDS and the Bayes’ error rate. We then establish the consistency of the sample estimate $T_{j}$ by showing that the ranking induced by the $T_{j}$s converges to that of the $\Delta_{j}$s under mild regularity conditions. We begin by providing the theorem below, which states the necessary conditions under which $\Delta_{j}$ can identify whether a subspace is aligned with the CDS.
Theorem 1.
Let $\mathbf{X}\in\mathbb{R}^{p}$ and $Y\in\{1,2\}$ follow model (1), $\boldsymbol{\Sigma}=\sum_{h=1}^{2}\pi_{h}\boldsymbol{\Sigma_{h}}$, and $\Delta_{j}$ be as defined in (11). If $\Delta_{j}\neq 0$, then $\boldsymbol{\Sigma}\mathbf{v}_{j}\not\perp\mathcal{S}_{D(Y\mid\mathbf{X})}$.
Theorem 1 is stated in terms of $\boldsymbol{\Sigma}\mathbf{v}_{j}$ rather than for any arbitrary $\mathbf{v}_{j}$ itself because, under model (1), the CDS for both LDA and QDA is characterized by the span of covariance-adjusted mean difference vectors of the form $\boldsymbol{\Sigma}^{-1}(\boldsymbol{\mu}_{h}-\boldsymbol{\mu}_{1})$ in $\mathcal{B}$. Thus, to determine whether a candidate direction is aligned with the CDS, we naturally assess $\boldsymbol{\Sigma}\mathbf{v}_{j}$, which places the direction in the same covariance-adjusted subspace as $\mathcal{S}_{D(Y\mid X)}$. Moreover, in the $\text{GEV}(\mathbf{M},\mathbf{N})$ formulation for an SDR basis when $\mathbf{N}=\boldsymbol{\Sigma}$, which is often the case, the relevant population subspace is likewise expressed through $\boldsymbol{\Sigma}\mathbf{v}_{j}$.
Thus, Theorem 1 establishes that $\Delta_{j}$ provides a criterion for determining whether the covariance-adjusted subspace spanned by a candidate direction $\mathbf{v}_{j}$ at least partially recovers $\mathcal{S}_{D(Y\mid X)}$. That is, $\Delta_{j}$ yields a measure for detecting whether any given subspace carries discriminatory signal. However, in general, many of the $p$ candidate directions produced by an SDR method may at least partially estimate the CDS. Therefore, we further establish $\Delta_{j}$ as a subspace criterion by demonstrating in the theorem below that any candidate direction with a larger $\Delta_{j}$ will yield a smaller Bayes’ error rate.
Theorem 2.
Let $\mathbf{X}\in\mathbb{R}^{p}$ and $Y\in\{1,2\}$ follow model (1). Let $\mathbf{v}_{j},\mathbf{v}_{\ell}$ exist such that $\mathbf{v}_{j}\neq\mathbf{v}_{\ell}$. Let $\varphi_{j}$ and $\Delta_{j}$ be as defined in (9) and (11), respectively. Then, we have the following results.
  1. 1.
    Let $\boldsymbol{\Sigma}_{1}=\boldsymbol{\Sigma}_{2}$. Then, $\Delta_{j}>\Delta_{\ell}$ if and only if $\varphi_{j}<\varphi_{\ell}$.
  2. 2.
    Let $\boldsymbol{\Sigma}_{1}\neq\boldsymbol{\Sigma}_{2}$, and suppose $\mathbf{v}_{j}^{\top}\boldsymbol{\Sigma}_{2}\mathbf{v}_{j}=\mathbf{v}_{\ell}^{\top}\boldsymbol{\Sigma}_{2}\mathbf{v}_{\ell}$. Then, $\Delta_{j}>\Delta_{\ell}$ if and only if $\varphi_{j}<\varphi_{\ell}$.
Although $\Delta_{j}$ allows $\boldsymbol{\Sigma}_{1}\neq\boldsymbol{\Sigma}_{2}$, it does not directly measure heteroscedasticity. Thus, for Theorem 2 under the QDA model, we impose that $\mathbf{v}_{j}^{\top}\boldsymbol{\Sigma}_{2}\mathbf{v}_{j}=\mathbf{v}_{\ell}^{\top}\boldsymbol{\Sigma}_{2}\mathbf{v}_{\ell}$. That is, for one population only, the variability in the subspaces spanned by $\mathbf{v}_{j}$ and $\mathbf{v}_{\ell}$ is similar. Under this imposed constraint, $\varphi_{j}$ is a monotonic function of $\Delta_{j}$. This assumption is easily satisfied, but not limited to, when either population covariance matrix is spherical. No additional assumptions are required under the LDA model.
By Theorem 2, we can directly compare the discriminatory strength of two subspaces using $\Delta_{j}$ under both the QDA and LDA decision rules. This result formalizes a key implication: Among a set of subspaces, the direction $\mathbf{v}_{j}$ with the largest $\Delta_{j}$ achieves the local minimum Bayes’ error rate. When $H=2$ under LDA, Lemma 2 yields $d=1$. Thus, we have that selecting the top-ranked direction by $\Delta_{j}$ directly estimates $\mathcal{S}_{D(Y\mid\mathbf{X})}$ and yields the local minimum Bayes’ error rate. Under QDA, Lemma 1 yields $d=\text{rank}(\mathcal{L}\cup\mathcal{Q})\geq 1$. Thus, while ordering by $\Delta_{j}$ still identifies directions with minimal 1D Bayes’ error rates, the $d$-dimensional error rate cannot be determined from individual rankings due to the lack of a tractable OER when $p>1$. However, as shown in the simulation studies in Section 5, combining top-ranked directions consistently yields substantial reductions in the estimated Bayes’ error rates, which demonstrates the practical gain of reordering subspaces under QDA. Therefore, we can reorder the subspaces resulting from an SDR method by their corresponding $\Delta_{j}$ values to maximize population separation.
The previous theoretical results establish $\Delta_{j}$ as an importance measure for evaluating and ordering directions in a DRS. Now we establish the consistency of its sample analogue $T_{j}$. Our interest is not solely in the pointwise convergence of each statistic but rather in the behavior of the entire vector $(T_{1},\ldots,T_{p})^{\top}$ and its properties as a ranking criterion. Specifically, for any vector $\boldsymbol{\theta}=(\theta_{1},\ldots,\theta_{p})^{\top}\in\mathbb{R}^{p}$ and $j\in\{1,\ldots,p\}$, we define the rank-order vector as
$\mathcal{R}(\boldsymbol{\theta})\coloneqq(r_{1}(\boldsymbol{\theta}),\ldots,r_{p}(\boldsymbol{\theta}))^{\top},\text{ where }r_{j}\coloneqq\sum_{i=1}^{p}\mathbf{1}\{\theta_{i}>\theta_{j}\}+1.$ (13)
Here, $r_{j}(\boldsymbol{\theta})$ denotes the relative rank-order of the $j^{\text{th}}$ component of $\boldsymbol{\theta}$, which assigns the lowest value of 1 to the largest entry. We first provide the necessary conditions under which uniform convergence of a finite-dimensional vector of estimators guarantees consistency of the induced rank-order in the lemma below.
Lemma 4.
Let $\boldsymbol{\theta}\coloneqq(\theta_{1},\ldots,\theta_{p})^{\top}\in\mathbb{R}^{p}$ be fixed, and let $\hat{\boldsymbol{\theta}}\coloneqq(\hat{\theta}_{1}^{(n)},\ldots,\hat{\theta}_{p}^{(n)})^{\top}\in\mathbb{R}^{p},n\in\mathbb{N}$. Suppose that $\hat{\theta}_{j}^{(n)}\xrightarrow{P}\theta_{j},\ j=1,\ldots,p$, and there exist an $\varepsilon>0$ such that $\left|\ \theta_{i}-\theta_{j}\,\right|\geq\varepsilon,\text{ for all }i\neq j$. Let $\mathcal{R}(\cdot)$ be as defined in (13). Then, $\mathbb{P}\left(\mathcal{R}(\hat{\boldsymbol{\theta}})=\mathcal{R}(\boldsymbol{\theta})\right)\to 1$.
Fundamental to our results in the theorems that follow, this lemma applies beyond the specific case of consistency for subspace ordering induced by the $T_{j}$s. The result requires that the entries of the population vector $\boldsymbol{\theta}$ be unique. We note that this condition is not restrictive because most ranking procedures assume distinct entries or break ties arbitrarily. Moreover, in our context, $\Delta_{j}$ is a continuous function of the model parameters and $\Delta_{j}=0$ if either $\boldsymbol{\mu}_{1}=\boldsymbol{\mu}_{2}$ (i.e., no signal exists) or $\mathbf{v}_{j}\perp\mathcal{S}_{D(Y\mid\mathbf{X})}$. To establish the rank-order consistency of the $T_{j}$s, we work under relaxed conditions with no assumption of multivariate normality. We impose only the following conditions.
Conditions.
  1. (C1)
    We have $\mathbf{X}_{hi}=\boldsymbol{\mu}_{h}+\boldsymbol{\varepsilon}_{hi}$ for $h=1,\ldots,H$ and $i=1,\ldots,n_{h}$.
  2. (C2)
    The vectors $\boldsymbol{\varepsilon}_{hi}$ are IID within each population $h$ such that $\mathbb{E}(\boldsymbol{\varepsilon}_{hi})=\mathbf{0}$ and $\text{Cov}(\boldsymbol{\varepsilon}_{hi})=\boldsymbol{\Sigma}_{h}\in\mathbb{S}_{+}^{p}$. Additionally, $\boldsymbol{\varepsilon}_{h\cdot}$ are independent across populations.
  3. (C3)
    Each component of $\boldsymbol{\varepsilon}_{hi}$ has a finite second moment.
Theorem 3.
Suppose conditions C1 - C3 hold, and let $\Delta_{j}$ and $T_{j}$, $j=1,\ldots,p$, be as defined in (11) and (12), respectively. Let $\mathcal{R}(\cdot)$ be defined as in (13). Then, we have $\mathbb{P}\left(\mathcal{R}(T_{1},\ldots,T_{p})=\mathcal{R}(\Delta_{1},\ldots,\Delta_{p})\right)\to 1$.
Theorem 3 formalizes the rank-order consistency of the sample-based $T_{j}$s with respect to the population-level $\Delta_{j}$s. This result relies on the fact that the estimated eigenvectors are root-$n$ consistent, as discussed in Section 2.3. This theorem ensures that, asymptotically, the ordering of subspaces by their empirical $T_{j}$ values recovers the population ordering induced by the $\Delta_{j}$s. Therefore, Theorems 2 and 3 establish that $T_{j}$ provides a consistent measure for evaluating and ordering directions in a DRS.
4 Subspace Criterion for a Categorical or Continuous Response
In the previous section, we established the use of $T_{j}$ as a criterion for ordering subspaces when the response is binary. When the response is categorical with $H\geq 2$, we extend this subspace ordering criterion by proposing an $F$-statistic, introduced by Fisher, (1925), within a DRS as an importance measure for subspaces, which we define as
$F_{j}\coloneqq\frac{\sum_{h=1}^{H}\widehat{\pi}_{h}\left(\bar{x}^{\ast}_{hj}-\sum_{i=1}^{H}\widehat{\pi}_{i}\bar{x}^{\ast}_{ij}\right)^{2}}{\sum_{h=1}^{H}\widehat{\pi}_{h}s_{hj}^{2\ast}},\quad j=1,\ldots,p.$ (14)
Moreover, the novelty of the $F_{j}$ measure lies in its ability to unify categorical and continuous responses under a single subspace criterion. Under the slicing-based framework in SDR, we can use the $H$ slices of $Y$ as categories in which the $F_{j}$ criterion can also be applied to continuous responses. Thus, for either a categorical or a continuous response, the $F_{j}$ criterion provides a measure of between-slice separation relative to within-slice variability for each subspace. For a categorical response, this criterion identifies subspaces that maximize overall population separation whereas, for a continuous response, ordering by $F_{j}$ yields subspaces that maximize slice separation, with the goal of identifying directions that preserve the conditional mean structure as the number of slices increases.
When the response $Y$ is discretized, either naturally through predefined populations or artificially through slicing, such that $\tilde{Y}\in\{1,\ldots,H\}$ with $H\geq 2$, we formalize $F_{j}$ as a subspace criterion by studying the theoretical properties of its population analogue,
$\Psi_{j}\coloneqq\frac{\displaystyle\sum_{h=1}^{H}\pi_{h}\left(\mathbf{v}_{j}^{\top}\boldsymbol{\mu}_{h}-\sum_{i=1}^{H}\pi_{i}\,\mathbf{v}_{j}^{\top}\boldsymbol{\mu}_{i}\right)^{2}}{\displaystyle\sum_{h=1}^{H}\pi_{h}\,\mathbf{v}_{j}^{\top}\boldsymbol{\Sigma}_{h}\mathbf{v}_{j}},\quad j=1,\ldots,p.$ (15)
If $H=2$, then Corollary 1 below formalizes the natural extension from $T_{j}$ to $F_{j}$ by establishing that $\Psi_{j}$ yields an identical ordering of subspaces as $\Delta_{j}$ and, as a result, retains the same theoretical properties for discrimination as $\Delta_{j}$.
Corollary 1.
Let H = 2. For each $\mathbf{v}_{j}$, $j=1,\ldots,p$, let $\Delta_{j}$ and $\Psi_{j}$ be as defined in (11) and (15), respectively. Let $\mathcal{R}(\cdot)$ be as defined in (13). Then, $\Delta_{j}\propto\sqrt{\Psi}_{j}$ and $\mathcal{R}(\Psi_{1},\ldots,\Psi_{p})=\mathcal{R}(\Delta_{1},\ldots,\Delta_{p})$.
In Theorem 1, we established that when the response is binary, $\Delta_{j}$ identifies directions that are at least partially aligned with $\mathcal{S}_{D(Y\mid\mathbf{X})}$. The following theorem shows that $\Psi_{j}$ retains this property in the multiclass setting and that for a continuous response we obtain a similar result in terms of $\mathcal{S}_{Y|\mathbf{X}}$.
Theorem 4.
Let $\mathbf{X}\in\mathbb{R}^{p}$ and $\tilde{Y}\in\{1,\ldots,H\}$. Let $\boldsymbol{\Sigma}=\sum_{h=1}^{H}\pi_{h}\boldsymbol{\Sigma_{h}}$ and $\Psi_{j}$ be as defined in (15). We then have the following results.
  1. 1.
    Suppose $\mathbb{E}[\mathbf{X}|\boldsymbol{\beta}^{\top}\mathbf{X}]$ is a linear function of $\boldsymbol{\beta}^{\top}\mathbf{X}$ such that $\text{span}(\boldsymbol{\beta})=\mathcal{S}_{Y|\mathbf{X}}$ and $\boldsymbol{\beta}\in\mathbb{R}^{p\times d}$. If $\Psi_{j}\neq 0$, then $\boldsymbol{\Sigma}\mathbf{v}_{j}\not\perp\mathcal{S}_{Y\mid\mathbf{X}}$.
  2. 2.
    Suppose $\mathbf{X}|\tilde{Y}$ follows model (1). If $\Psi_{j}\neq 0$, then $\boldsymbol{\Sigma}\mathbf{v}_{j}\not\perp\mathcal{S}_{D(Y\mid\mathbf{X})}$.
Theorem 4 establishes that $\Psi_{j}$ provides a criterion for detecting informative subspaces that guarantees partial alignment with $\mathcal{S}_{Y|\mathbf{X}}$ under the linearity condition or with $\mathcal{S}_{D(Y|\mathbf{X})}$ under model (1). For the continuous response case, the linearity condition is a mild and standard assumption in SDR that is also satisfied whenever $\mathbf{X}$ follows an elliptical distribution. To extend these results to the sample setting, we establish that $F_{j}$ yields a consistent ranking of directions relative to $\Psi_{j}$, as formalized in Theorem 5.
Theorem 5.
Suppose conditions C1 - C3 hold, and let $F_{j}$ and $\Psi_{j}$, $j=1,\ldots,p$, be as defined in (14) and (15), respectively. Let $\mathcal{R}(\cdot)$ be defined as in (13). Then, we have $\mathbb{P}\left(\mathcal{R}(F_{1},\ldots,F_{p})=\mathcal{R}(\Psi_{1},\ldots,\Psi_{p})\right)\to 1$.
Theorems 4 and 5 establish $F_{j}$ as a consistent sample-based subspace criterion that preserves the ranking induced by $\Psi_{j}$, thereby providing a measure for evaluating the relevance of candidate directions. We note that, in contrast to the binary case where $\Delta_{j}$ yields a direct monotonic relationship with the Bayes error rate, such a result is not available for $\Psi_{j}$ because no general tractable OER for LDA or QDA exists when $H>2$. However, our simulation results and real-data applications demonstrate the practical effectiveness of ordering subspaces by $F_{j}$, which identifies directions with greater overall class separation and, as a result, often significantly reduces the estimated Bayes’ error rate.
For a continuous response, a similar limitation arises because we cannot evaluate any population-level measure of optimality against $\Psi_{j}$. However, in the sample case, when $\text{span}(\boldsymbol{\beta})=\mathcal{S}_{Y|\mathbf{X}}$ must be estimated by $\widehat{\boldsymbol{\beta}}$, which corresponds to the SDR method-specific eigenvectors $\widehat{\mathbf{V}}$, the distance between the estimated and true subspaces can be quantified by
$\mathcal{D}(\mathcal{S}_{\boldsymbol{\beta}},\mathcal{S}_{\hat{\boldsymbol{\beta}}})=\frac{{\|\mathbf{P}_{\boldsymbol{\beta}}-\mathbf{P}_{\hat{\boldsymbol{\beta}}}\|}_{F}}{\sqrt{2d}}.$ (16)
[…]
5 Simulation Studies
5.1 Simulation for Binary Response
We used Monte Carlo (MC) simulations to demonstrate the efficacy of reordering the DRS of an SDR method using the $T_{j}$ criterion contrasted to the eigenvalue magnitude. The PCA, SAVE, SIR-II, and SSDR methods were implemented as described in Section 2.3. Note that we implemented the SIR-II method rather than the classical SIR method because SIR relies only on first-order moments. This fact makes its eigenvalues proportional to $T_{j}$ and thus yields the same ordering. Using the $F_{j}$ criterion, we provide additional simulations in the Supplementary Material for categorical responses with $H>2$. We refer to the eigenvalue-ordered SDR methods by their standard names and the $T_{j}$-ordered versions as $PCA_{T}$, $SAVE_{T}$, $SIR\text{-}II_{T}$, and $SSDR_{T}$, respectively.
We considered parameter configurations from multivariate normal populations with $p=50$ for the MC simulation study. For configurations Q1-Q3 below, we used QDA as the supervised classifier.
  • Configuration Q1: $\boldsymbol{\mu}_{1}=\mathbf{0}_{p}$, $\boldsymbol{\mu}_{2}$ is a $p\times 1$ vector with IID entries from the $\mathcal{N}(0,1)$ distribution, $\boldsymbol{\Sigma}_{1}=\mathbf{I}_{p}$, $\boldsymbol{\Sigma}_{2}=\begin{bmatrix}\phantom{-}3&-2\\ -2&\phantom{-}3\end{bmatrix}\oplus\mathbf{I}_{p-2}$, and $d=2$.
  • Configuration Q2: $\boldsymbol{\mu}_{1}=\mathbf{0}_{p}$, $\boldsymbol{\mu}_{2}=(\mathbf{1}_{5},-\mathbf{1}_{5},\mathbf{0}_{p-10})$, $\boldsymbol{\Sigma}_{1}=\mathbf{I}_{p}$, and $\boldsymbol{\Sigma}_{2}=[\rho\mathbf{I}_{b}+(1-\rho)(\mathbf{J}_{b}-\mathbf{I}_{b})]\oplus\mathbf{I}_{p-b}$, where $\mathbf{J}_{p}$ is a $p\times p$ matrix of ones, $b=5$, $\rho=0.99$, and $d=6$.
  • Configuration Q3: $\boldsymbol{\mu}_{1}=\mathbf{0}_{p}$, $\boldsymbol{\mu}_{2}=(\mathbf{0}_{p-1},1)^{\top}$, $\boldsymbol{\Sigma}_{1}=\boldsymbol{\Sigma}_{2}=(\sigma_{ij})$, where $\sigma_{ii}=2,\text{ for all }i\neq p$, $\sigma_{pp}=1$, $\sigma_{ij}=0,\text{ for all }i\neq j$, and $d=1$.
We used LDA as the supervised classifier for configurations L1-L3 below.
  • Configuration L1: $\boldsymbol{\mu}_{1}=\mathbf{0}_{p}$, $\boldsymbol{\mu}_{2}=\mathbf{1}_{p}$, $\boldsymbol{\Sigma}_{1}=\boldsymbol{\Sigma}_{2}=(1-\rho)\mathbf{I}_{p}+\rho\mathbf{J}_{p}$, $\rho=0.25$, and $d=1$.
  • Configuration L2: $\boldsymbol{\mu}_{1}=(1,1,\mathbf{0}_{p-2})^{\top}$, $\boldsymbol{\mu}_{2}=-\boldsymbol{\mu}_{1}$, and $\boldsymbol{\Sigma}_{1}=\boldsymbol{\Sigma}_{2}=\mathbf{I}_{2}\oplus s^{2}[(1-\rho)\mathbf{I}_{p-2}+\rho\mathbf{J}_{p-2})]$, where $s^{2}=10$, $\rho=0.99$, and $d=1$.
  • Configuration L3: The same setting as configuration Q2 except $b=20$. Thus, $d=20$.
Figure 2: Estimated conditional error rate (CER^)($\widehat${CER}) plots for the simulation described in Section 5.1 for contrasting eigenvalue-ordered SDR methods (labeled by their standard names) with their $T_{j}$-ordered counterparts (denoted by the superscript T). All SDR methods used the QDA classifier, and the horizontal line represents the median $\widehat{CER}$ using QDA without dimension reduction.
For each configuration, we generated 5,000 observations from each population. We varied the training sample sizes to simulate ill-conditioned to well-conditioned estimation of the discriminant parameters. The class-specific training sample sizes for the QDA configurations were $n_{h}=p+1,\ 2p,\ \text{and }5p,\ h\in\{1,2\}$. For the LDA configurations, the training sample sizes were n=p+1, 2p, 5pn=p+1,\ 2p,\ 5p with equal prior class probabilities. We used the remaining observations as the test set. For each SDR method, we projected the training and test sets from $p$ to $d=\text{dim}(\mathcal{S}_{D(Y|\mathbf{X})})$ dimensions, then applied the respective classifier and recorded the estimated conditional error rate, denoted by $\widehat{CER}$. We also recorded the $\widehat{CER}$ of the full-feature data using the respective classifier without dimension reduction. This process was replicated 1,000 times for each configuration. We summarized the MC simulation results in Figures 2 and 3 for the QDA and LDA configurations, respectively, by displaying the distributions of the $\widehat{CER}$s for each SDR method. The horizontal line denotes the median $\widehat{CER}$ of the full-dimensional classifier with no dimension reduction.
From Figure 2, we clearly see that, compared to ordering by eigenvalue magnitude, ordering the DRS by $T_{j}$ can significantly reduce the $\widehat{CER}$. In most cases, we found that eigenvalue ordering often yielded SDR methods with $\widehat{CER}$ values much larger than those of the full-feature QDA, whereas ordering by $T_{j}$ resulted in substantial improvements relative to the full-feature $\widehat{CER}$s. The SIR-II method exhibited the largest gain when it was reordered by the $T_{j}$ criterion. In fact, the $SIR\text{-}II_{T}$ method often achieved the minimum error rates. In contrast, the standard SIR-II consistently produced $\widehat{CER}$s around 0.50. When $n_{h}=p+1=51$ for configuration Q1, the median $\widehat{CER}$ achieved by the full-feature QDA was 0.4287. However, regardless of sample size, the $SIR\text{-}II_{T}$ method yielded an impressive median $\widehat{CER}$ of approximately 0.02. In larger-sample scenarios, the $T_{j}$-ordered SDR methods performed as well as or, in some cases, significantly better than the full-feature QDA. For instance, when $n_{h}=5p=250$ for configuration Q2, the $SAVE_{T}$, $SIR\text{-}II_{T}$, and $SSDR_{T}$ methods achieved the lowest error rates, with $SSDR_{T}$ performing best. Additionally, for configuration Q3 with $n_{h}=250$, the $\textit{PCA}_{T}$ method achieved the minimum median $\widehat{CER}$ of 0.3484, compared to 0.4472 for QDA and 0.4994 for standard PCA. In contrast, all other SDR methods yielded equal or higher error rates than the full-dimensional QDA.
Figure 3: Estimated conditional error rate (CER^)($\widehat${CER}) plots for the simulation described in Section 5.1, contrasting eigenvalue-ordered SDR methods (labeled by their standard names) with their $T_{j}$-ordered counterparts (denoted by the superscript $T$). LDA was used as the supervised classifier for all SDR methods. The median $\widehat{CER}$ using LDA without dimension reduction is represented by the horizontal line.
Under LDA from Figure 3, the $T_{j}$-ordered SDR methods clearly achieved superior performance compared to their eigenvalue-ordered counterparts. In configuration L1, PCA and $PCA_{T}$ performed similarly. However, in configuration L2, PCA produced a consistent median $\widehat{CER}$ of approximately 0.50. For $n=p+1=51$ and $n=2p=100$, $PCA_{T}$ yielded the minimum error rates with median $\widehat{CER}$s around 0.0950 compared to the median $\widehat{CER}$s for LDA of 0.1820 and 0.1235, respectively. Under Model 1, when the parameters are known, LDA yields the Bayes optimal subspace. Thus, for large samples, the full-feature LDA achieved equal or better performance than LDA did after we used the SDR methods to reduce the feature space, which was expected. However, in the large-sample scenarios, ordering subspaces by $T_{j}$ resulted in performance nearly identical to the optimal full-feature LDA whereas eigenvalue ordering produced significantly larger $\widehat{CER}$s for certain methods across the configurations.
5.2 Simulation for Continuous Response
Through MC simulations, we illustrated the efficacy of reordering the DRS by using the $F_{j}$ criterion when the response was continuous. We used the same SDR methods as in Section 5.1 and, likewise, referred to the $F_{j}$-ordered SDR methods as $PCA_{F}$, $SAVE_{F}$, $SIR\text{-}II_{F}$, and $SSDR_{F}$. The eigenvalue-ordered SDR methods are referred to by their original names. For all SDR methods, we set the number of slices to $H=5$.
Similar to our simulation set up in Section 5.1, we set $p=50$ for all parameter configurations. We varied the three training sample sizes of $n=H\cdot(p+1)$, $H\cdot(2p)$, and $H\cdot(5p)$ to again simulate poorly-posed to well-conditioned estimation of the conditional moments within each slice. For each SDR method, to evaluate subspace estimation accuracy and, hence, predictor–response preservation, we recorded the estimated subspace distance $\mathcal{D}$ in (16) between the true basis $\boldsymbol{\beta}\in\mathbb{R}^{p\times d}$ and the estimated basis $\widehat{\boldsymbol{\beta}}\in\mathbb{R}^{p\times d}$. We replicated this process 1,000 times for each parameter configuration and summarized the results in Figure 4, which displays the distribution of $\mathcal{D}$ for each SDR method.
The configurations used are similar to those in Zeng et al., (2024). For configurations D1 and D2 below, $\mathbf{X}_{i}$ follows a multivariate normal distribution with $\boldsymbol{\mu}=\mathbf{0}_{p}$ and $\boldsymbol{\Sigma}=AR(0.50)$, where $AR(0.50)$ is a $p\times p$ auto-regressive matrix whose $(i,j)^{\text{th}}$ element is $0.50^{|i-j|}$. For each model, $\varepsilon_{i}$ follows the $\mathcal{N}(0,1)$ distribution. The configurations are as follows:
  • Configuration D1: $Y_{i}=\boldsymbol{\beta}^{\top}\mathbf{X}_{i}+\varepsilon_{i}$, where $\boldsymbol{\beta}\in\mathbb{R}^{p}$ with IID entries from $\mathcal{N}(0,1)$.
  • Configuration D2: $Y_{i}=(\boldsymbol{\beta}_{1}^{\top}\mathbf{X}_{i})\cdot\exp\left(\boldsymbol{\beta}_{2}^{\top}\mathbf{X}_{i}+\varepsilon_{i}\right)$, where $\beta_{1i}$’s and $\beta_{1j}$’s have IID entries from the uniform(0.30, 0.60) distribution for $1\leq i\leq 30$ and $1\leq j\leq 15$, $\beta_{1j}$’s have IID entries from the uniform(-0.30, -0.60) for $16\leq j\leq 30$, and $\beta_{ij}=0$ otherwise.
  • Configuration D3: The same model as configuration D2, except $\mathbf{X}_{i}$ has a non-elliptical distribution such that $\mathbf{X}_{i}\sim 0.40\mathcal{N}(\boldsymbol{\mu}_{1},\boldsymbol{\Sigma}_{1})+0.20\mathcal{N}(\boldsymbol{\mu}_{2},\boldsymbol{\Sigma}_{2})+0.40\mathcal{N}(\boldsymbol{\mu}_{3},\boldsymbol{\Sigma}_{3})$, where $\boldsymbol{\mu}_{1}=(-\mathbf{1}_{30},\mathbf{0}_{p-30})$, $\boldsymbol{\Sigma}_{1}=AR(0.10)$, $\boldsymbol{\mu}_{2}=\mathbf{0}_{p}$, $\boldsymbol{\Sigma}_{2}=AR(0.50)$, $\boldsymbol{\mu}_{3}=-\boldsymbol{\mu}_{1}$, and $\boldsymbol{\Sigma}_{3}=AR(0.90)$.
Figure 4: Estimated subspace distance $\mathcal{D}$, given in (16), plots for contrasting eigenvalue-ordered SDR methods (labeled by their standard names) with their $F_{j}$-ordered counterparts (denoted by the superscript $F$). The number of slices was set to $H=5$ for all SDR methods.
From Figure 4, we found that the $F_{j}$-ordered SDR methods achieved superior subspace estimation compared to the eigenvalue-ordered counterparts. For every SDR method, configuration, and sample size, the $F_{j}$-ordering resulted in either comparable performance to eigenvalue ordering or, in most cases, significantly lower $\mathcal{D}$ values. For configuration D1, which was a single-index model with $d=1$, PCA resulted in a median $\mathcal{D}$ of 1.00, regardless of sample size, which indicated that no predictor–response information was preserved. This result is not surprising because PCA considers only $\mathbf{X}$ and is an unsupervised SDR method. In contrast, for $n=H\cdot(p+1)=55$, $H\cdot(2p)=500$, and $H\cdot(5p)=1{,}250$, $PCA_{F}$ achieved median $\mathcal{D}$ values of 0.3206, 0.2303, and 0.1692, respectively. Moreover, for $n=55$ and 500, $PCA_{F}$ yielded the minimum $\mathcal{D}$ values when contrasted to all other methods. This difference illustrated that PCA, traditionally an unsupervised method, can become a pseudo-supervised method when it incorporates the $F_{j}$-ordering criterion, thus enabling it to compete with or outperform traditional supervised SDR methods.
For all configurations, the SAVE and SDRS methods performed poorly for small sample sizes whereas the $SAVE_{F}$ and $SDRS_{F}$ methods yielded modest to substantial improvements. With larger sample sizes, the performances of SAVE and SDRS were similar across orderings with modest advantages for the $F_{j}$-based variants. Configurations D2 and D3 were multiple-index models with $d=2$. Here, similar to PCA, the SIR-II method yielded no subspace recovery when ordered by eigenvalue magnitude whereas the $SIR\text{-}II_{F}$ method achieved superior predictor–response preservation, often yielding the minimum $\mathcal{D}$ values.
6 Real Data Applications
In this section, we present an application of our proposed $T_{j}$ subspace criterion to high-dimensional gene expression data and an application of our $F_{j}$ criterion to continuous response data for housing price prediction. For simplicity of presentation, we considered two-fold validation for both data applications. Additional real-data applications are provided in the Supplementary Material, including several repeated 10-fold cross-validation error rate comparisons for each SDR method using both $T_{j}$ and $F_{j}$, similar to those in Section 5.1. Moreover, we provide results for the DR method within our proposed subspace ordering framework. We also include an analysis of brain cancer data illustrating the relationship between the empirical distribution of $T_{j}$ and the empirical error rate distribution.
6.1 Application to High-Dimensional Gene Expression Data
Here, we present our analysis of high-dimensional gene expression data for two leukemia subtypes from Golub et al., (1999). The data were obtained from bone marrow samples of 72 patients: 47 diagnosed with acute lymphoblastic leukemia (ALL) and 25 with acute myeloid leukemia (AML). Affymetrix Hgu6800 chips were used to extract 7,129 gene expression levels from each patient.
Figure 5: Leukemia data discussed in Section 6.1 in the estimated SIR-II subspaces corresponding to the largest eigenvalue (left plot) and the largest $T_{j}$ value (right plot). QDA was used for the estimated decision boundaries. Vertical jitter was added for visualization.
The authors provided training and testing sets consisting of 38 and 34 patients, respectively. For the training set, we estimated the DRS using the SIR-II method. To ensure the generalized eigenvalue problem was solvable, we applied Tikhonov regularization to the sample covariance matrix such that $\widehat{\mathbf{N}}=\mathbf{S}+\gamma\mathbf{I}_{p}$ with $\gamma\coloneqq 10^{-6}$. We obtained the $SIR\text{-}II_{T}$ DRS by reordering the SIR-II basis using the $T_{j}$ criterion. The dimensionality of the training and testing data was reduced to $d=1$, which we determined via validation using the training set. Next, we estimated the QDA decision boundary using the reduced training data in both DRSs. Our results are shown in Figure 5, which displays the reduced testing data in the estimated SIR-II and $SIR\text{-}II_{T}$ subspaces, respectively, along with the corresponding QDA decision boundaries derived from the reduced training data.
From Figure 5, we readily see that the SIR-II subspace associated with the largest eigenvalue yielded no discriminatory information. That is, no response information was preserved, and all patients were classified as ALL by QDA. This result yielded a $\widehat{CER}=0.4117$. In contrast, the SIR-II subspace associated with the largest $T_{j}$ achieved clear separation between the ALL and AML samples, with a substantially lower $\widehat{CER}=0.1471$. Every patient with ALL was correctly identified, and only five patients with AML were misclassified. Similar results were obtained under LDA, where SIR-II performed comparably, and $SIR\text{-}II_{T}$ yielded a slightly higher $\widehat{CER}=0.1764$.
We attributed the superior performance of the $SIR\text{-}II_{T}$ method to the $T_{j}$ criterion, which provided a data-driven measure that incorporates predictor–response information and yielded a more stable ordering criterion than eigenvalue magnitudes did. In particular, due to singularity issues, the largest eigenvalue was $10^{4}$ and non-unique across the first 7,092 eigenvectors. In contrast, the $T_{j}$ values exhibited clear separation, where the largest value was 5.41, and the second largest was 1.44. The remaining $T_{j}$ values decreased gradually and were unique across all subsequent subspaces. Figures showing the eigenvalue and $T_{j}$ orderings are provided in the Supplementary Material. The eigenvector associated with the largest eigenvalue yielded $T_{1}=0.01$, ranking 6,968th among all $T_{j}$ values whereas the subspace that achieved the largest $T_{j}=5.41$ corresponded to the smallest eigenvalue of 0.0001. Thus, under the traditional eigenvalue-ordering framework, the most informative subspace would not have been selected; however, using the $T_{j}$ criterion, we identified this subspace as the most informative and stable direction.
6.2 Application to Real Data with a Continuous Response
In this section, we analyzed the Boston housing price data, originally compiled from the 1970 U.S. Census. The dataset contained 506 observations, each representing a census tract within the Boston metropolitan area. For each tract, various socioeconomic, environmental, and structural characteristics were recorded, such as the per capita crime rate by town, average number of rooms per dwelling, full-value property tax rate per \$10,000, and others. The response variable was the median value of owner-occupied homes per \$1,000. See Harrison and Rubinfeld, (1978) for additional details.
Figure 6: Boston housing data discussed in Section 6.2 in the estimated SAVE subspaces corresponding to the largest eigenvalue (left plot) and the largest $F_{j}$ value (right plot).
We randomly partitioned the data into training (80%) and testing (20%) subsets for model estimation and evaluation. As in Section 6.1, we estimated the DRS using the training data for both the SAVE and $SAVE_{F}$ methods, thereby reducing the data dimensionality to $d=1$, which we determined via the training data. Figure 6 presents our results, displaying scatter plots between the median housing prices for the testing set and the reduced testing data in the estimated SAVE and $SAVE_{F}$ subspaces, respectively. From Figure 6, we again found that the subspace associated with the largest eigenvalue captured no response information. In the traditional first SAVE subspace, the resulting testing mean squared error (MSE) from a simple linear regression model was 59.91. The model was clearly insignificant and yielded an MSE comparable to that obtained when we used only the training sample mean, which we found to be 62.82. In contrast, the $F_{j}$ criterion identified a subspace, specifically the fourth SAVE direction, that captured a meaningful predictor-response structure and reduced the testing MSE to 28.05 when it was fitted with a second-degree polynomial model.
This real-data application highlights an important limitation of ordering by eigenvalue magnitude. Although larger eigenvalues typically correspond to directions of greater variability and are often assumed to capture more information, this assumption generally does not hold in practice. In this case, the SAVE subspace associated with the largest eigenvalue did exhibit greater overall variance than the $SAVE_{F}$ subspace. However, this increased variance arose because the eigenvector corresponding to the largest eigenvalue was sparse and effectively projected the data onto two distinct spans. In contrast, the $F_{j}$ criterion considered meaningful variability relative to the response and predictors, which yielded a subspace that captured a more informative structure than just incidental variability.
7 Discussion
We proposed using criteria other than eigenvalue magnitude for ordering subspaces in SDR. Traditionally, researchers have used eigenvalues to determine the basis for a DRS; however, we have demonstrated that eigenvalues generally do not correspond to the predictive relevance of a subspace. Therefore, we proposed and established theoretical results for the $T_{j}$ and $F_{j}$ criteria, which rank subspaces by their relevance to the response. The proposed criteria provide a straightforward and interpretable alternative that aligns the subspace ordering step with the goal of SDR—namely, maximizing predictor-response preservation.
Although we do not claim that reordering subspaces by our proposed criteria universally yields the best predictive performance, we have found considerable evidence through both simulations and real-data applications that ordering a DRS by the $T_{j}$ and $F_{j}$ criteria generally results in lower misclassification rates and more accurate subspace estimation than ordering by eigenvalue magnitude. Specifically, in high-dimensional or ill-conditioned settings, eigenvalues can yield an unstable subspace criterion with non-unique values. Moreover, when the leading eigenvectors are sparse, the associated eigenvalues no longer reflect response-related variability but instead capture incidental variance. In contrast, our proposed criteria provide a stable and superior ordering of subspaces. This fact reflects that $T_{j}$ and $F_{j}$ serve as data-driven subspace criteria that, to some extent, correct for subspace estimation variability by emphasizing maximum population or slice separation. Additionally, ordering subspaces by either $T_{j}$ or $F_{j}$ enables unsupervised methods, such as PCA, to behave in a pseudo-supervised manner and, in some cases, compete with or even outperform conventional supervised SDR methods.
We view this work as an initial step toward developing a broader class of prediction-specific subspace criteria. Our goal is to have demonstrated through the proposed $T_{j}$ and $F_{j}$ criteria that subspace ordering can and often should be informed by the objectives of the supervised learner rather than defaulting to eigenvalue magnitude. In this paper, we considered discrimination and regression settings; however, deriving alternative subspace criteria tailored to other learning objectives remains for future research. Moreover, we assumed that the structural dimension of the CS was known; jointly estimating the structural dimension while simultaneously ordering subspaces using a predictive importance measure is another related future research topic. Such extensions would allow SDR methods to integrate more tightly with the intended supervised model and produce a DRS that is not only statistically sufficient but also optimally informative for prediction.
Code Availability
All reproducible code for the simulations, real data applications, and graphical figures is available at https://github.com/DerikTBoonstra/Subspace_Ordering. Moreover, the $T_{j}$ and $F_{j}$ criteria, along with all of the SDR methods used, are implemented in the working sdr package available at https://github.com/DerikTBoonstra/sdr.
Acknowledgements
The authors received no financial support from any funding agency in the public, commercial, or not-for-profit sectors.